an electric fan is rotating at 3.0 rad/s when at time t=0 it begins to constantly accelerate at 2.00 rad/s^2. The initial linear acceleration of the edge of the spinning part of the fan has a magnitude of 1.30 m/s^2. What is the radius of the spinning part of the fan?


1
Expert's answer
2022-03-07T10:19:08-0500

The linear acceleration "a" is lincked with the angular one "\alpha" as follows:


"a = r\alpha"

where "r" is the radius of the circle. Thus, since "a= 1.3m/s^2, \alpha = 2rad/s^2", obtain:


"r = \dfrac{1.3}{2} = 0.650m"

Answer. 0.650m.